# Early-Exit and Instant Confidence Translation Quality Estimation

### v1

@@ 6 Conclusion
We introduced three approaches to improve the efficiency of quality estimation. Our instant confidence Early-Exit COMET achieves comparable performance to Monte Carlo dropout methods while drastically lowering computational overhead. Combining our model with a simple early-exit strategy, we can compute comparable quality estimation scores without having to compute the full quality estimation model. Finally, combining Early-Exit COMET with an upper confidence bound bandit, we speed up candidate reranking for machine translation by a factor of almost $2$ with negligible impact on translation quality.
@@
Reranking can also be combined with beam search to improve quality at generation time. Future work could apply our ideas to improve reranking efficiency for model generation. Moreover, in Appendix˜G we describe a prototype of Partial COMET, a quality estimation model that is able to robustly evaluate incomplete generations. This can be used to prune incomplete candidates, thus saving unnecessary computation of the very expensive generative model.
@@
Regarding the results in Table˜1, it is possible that there is a causal trade-off between the two correlation scores, the correlation with human scores and the correlation with the true error. For example, it could be easier to predict the confidence of a model that performs worse. However, making stronger claims would require a more thorough mathematical treatment, which is outside of the focus of this work.
@@ Acknowledgments
We thank the EAMT committee for sponsoring this research. We thank the Vector Stiftung for supporting Béni Egressy’s work. Part of this work received support from the European Union’s Horizon research and innovation programme under grant agreement No 101135798, project Meetween (My Personal AI Mediator for Virtual MEETtings BetWEEN People).
@@ Appendix A Model Architecture and Training Details
Table˜3 provides details of the model architecture and training hyperparameters that were used to train the COMET models in this paper.
[…]
@@
Regression head $(4{\times}768)\times 2048{\times}1024\times(1$ or $2)$
@@ Appendix C Additional Plots for Instant Confidence COMET
We provide additional plots showing the quality of the Instant Confidence COMET error predictions. In Figure˜6 we plot the average predicted error versus the true error to indicate average alignment. We also include correlation scores for selected layers showing the correlation between the predicted and true errors, e.g., $0.44$ for layer $9$ scores. In Figure˜7 we plot the average error for different confidence bins based on Instant Confidence. We see that the true error decreases as the predicted confidence increases. The plot also indicates that the score predictions with the highest and lowest true errors are reliably identified by the predicted instant confidence values.
@@ Appendix D Baseline Early Exit Algorithms
Algorithm˜3 exits the COMET evaluation after a fixed, predefined layer, exit layer $k$.
Algorithm˜4 exits the COMET evaluation when Early-Exit COMET scores from three consecutive layers are close to each other, or more precisely, have variance below a chosen threshold $\tau$.
@@
3: Compute $L_{i}(s,t)$ from $L_{i-1}(s,t)$ $\triangleright$ next layer
@@
3: Compute $L_{i}(s,t)$ from $L_{i-1}(s,t)$ $\triangleright$ next layer
[…]
5:if $\mathrm{Var}[\hat{y}_{i-2:i}]{<}\tau$ then return $\hat{y}_{i}$ $\triangleright$ early-exit
@@ Appendix E Upper Confidence Bound Bandit Ablations
We now describe two variations to the Upper Confidence Bound Bandit algorithm described in Section˜4: (1) starting at later layers, and (2) balancing exploration and exploitation.
@@
Given the significant jump in COMET score accuracy within the first few layers of the COMET model (Section˜3.1), we decided to explore initializing the algorithm with different starting layers. This carries a higher initial cost as we have to run the first few layers of the COMET model for all candidates but could lead to better-informed exploration with the remaining budget. The results in Section˜3.1 show that this leads to only marginal improvements.
@@
The heart of the multi-armed bandit problem is the exploration-exploitation trade-off. In our algorithm this trade-off is controlled by the hyperparameter $\gamma$. The higher one chooses $\gamma$, the higher the Upper Confidence Bound scores for uncertain candidates will be, and therefore the more likely the algorithm will be to explore many candidates. On the other hand, a low $\gamma$ will lead the algorithm to go deep with the most promising candidates, i.e., those with the highest estimated scores. We provide results for two different values for $\gamma$ in Section˜3.1, which shows that the default choice is likely the most apt.
@@
At the start (budget, $B=5\%$), only the first layer COMET scores are explored for (almost) all candidates, and by the end (budget, $B=100\%$) the full COMET scores are calculated for all candidates. However, the distributions in between reveal that the bandit often explores the candidates up until certain modal layers. These can be clearly seen (e.g. for budget $B=50\%$) to be layers $3$, 44, $11$, $18$, and $24$. Looking at Figure˜6, it can be seen that these values correspond to layers where the predicted error rates (uncertainties) drop significantly, meaning that the evaluation model is now more certain about these scores relative to scores for other candidates. This explains why the UCB bandit algorithm would then prefer to explore less certain candidates with higher UCB scores.
@@ Appendix F Reranking Results in Tabular Format
Table˜4 provides the results from Section˜3.1 in tabular format.
@@
Table 4: Quality of the candidates returned by the Upper Confidence Bound bandit. Quality is measured in terms of the average final candidate score and the proportion to top-1 candidates selected. We plot these measures for various evaluation budgets. Cost (or budget) is given relative to calculating the full COMET scores for all candidates ($100\%$). Visualized in Section˜3.1.
@@ Appendix G Partial COMET
In this section, we introduce Partial COMET (addressing non-standard input obstacle), ensuring reliable quality assessments for incomplete outputs. We show that also beyond evaluation, partial quality estimation can assess translation quality early in the translation generation process, and help discard unpromising candidates from beam search or sampling-based methods, allowing only the most promising candidates to be further expanded. (Appendix˜G)
@@ G.1 Modeling
This way, the quality estimation model predicts final translation scores based on just the translation prefix. See Table˜6.
@@ G.2 Results
We show the results, as measured by Pearson correlation across WMT23 in Table˜7. The full COMET strongly underperforms on partial generations. Therefore, this quality estimation model is not suitable for evaluating incomplete segments during generation. In contrast, the partial quality estimation model correlates much more. This can also be due to human annotators overfocusing on the beginning of the translations when providing assessment scores (Magooda and Marcjan, 2020), or the quality estimation system picking up spurious correlations that give away the score (Zouhar et al., 2024).
@@
We try all combinations of pruning bottom 0%, 25%, 50%, or 75% at each of 25%, 50%, and 75% target lengths. The results in Section˜3.1 show that by using Partial COMET, we are more likely to prune lower-quality candidates than with the original COMET, which itself performs better than random pruning.

### v2

@@ 7 Conclusion
We introduced three approaches to improve the efficiency of quality estimation. Our instant confidence Early-Exit COMET achieves comparable performance to Monte Carlo dropout methods while drastically lowering computational overhead. Combining our model with a simple early-exit strategy, we can compute comparable quality estimation scores without having to compute the full quality estimation model. Combining Early-Exit COMET with an upper confidence bound bandit, we speed up candidate reranking for machine translation by a factor of almost $2$ with negligible impact on translation quality. Finally, the confidence scores can also inform which translations to human-annotate.
@@
  • When human-annotating only a portion of translations, prioritize those with low Instant Confidence COMET confidence.
@@
Reranking can also be combined with beam search to improve quality at generation time. Future work could apply our ideas to improve reranking efficiency for model generation. Moreover, in Appendix G we describe a prototype of Partial COMET, a quality estimation model that is able to robustly evaluate incomplete generations. This can be used to prune incomplete candidates, thus saving unnecessary computation of the very expensive generative model.
@@
Regarding the results in Table 1, it is possible that there is a causal trade-off between the two correlation scores, the correlation with human scores and the correlation with the true error. For example, it could be easier to predict the confidence of a model that performs worse. However, making stronger claims would require a more thorough mathematical treatment, which is outside of the focus of this work.
@@ Acknowledgments
We thank the EAMT committee for sponsoring this research. We thank the Vector Stiftung for supporting Béni Egressy’s work. Part of this work received support from the European Union’s Horizon research and innovation programme under grant agreement No 101135798, project Meetween (My Personal AI Mediator for Virtual MEETtings BetWEEN People). This research has been funded in part by a Swiss National Science Foundation award (project 201009) and a Responsible AI grant by the Haslerstiftung.
@@ Appendix A Model Architecture and Training Details
Table 3 provides details of the model architecture and training hyperparameters that were used to train the COMET models in this paper.
Unless otherwise specified, the configuration of referenceless COMET is used. Our models are largely compatible with the upstream COMET repository and can be reproduced based on our code. We run all our experiments on Nvidia A100 (40GB) GPUs, taking about 8 hours to train a single model (2.2GB) for 5 epochs. For each setting, we train a single model and report its performance. In total, all experiments, including preliminary ones, amounted to approximately 20$\times$8 hours = 160 hours of compute on the aforementioned GPU. We base our experiments on a modified COMET v2.2.4 codebase with other package versions listed as dependencies in this version.
@@
[…]
@@ Appendix C Additional Plots for Instant Confidence COMET
We provide additional plots showing the quality of the Instant Confidence COMET error predictions. In Figure 7 we plot the average predicted error versus the true error to indicate average alignment. We also include correlation scores for selected layers showing the correlation between the predicted and true errors, e.g., $0.44$ for layer $9$ scores. In Figure 8 we plot the average error for different confidence bins based on Instant Confidence. We see that the true error decreases as the predicted confidence increases. The plot also indicates that the score predictions with the highest and lowest true errors are reliably identified by the predicted instant confidence values.
@@ Appendix D Baseline Early Exit Algorithms
Algorithm 3 exits the COMET evaluation after a fixed, predefined layer, exit layer $k$.
Algorithm 4 exits the COMET evaluation when Early-Exit COMET scores from three consecutive layers are close to each other, or more precisely, have variance below a chosen threshold $\tau$.
@@
3:   Compute $L_{i}(s,t)$ from $L_{i-1}(s,t)$triangleright next layer
@@
3:   Compute $L_{i}(s,t)$ from $L_{i-1}(s,t)$triangleright next layer
4:   $\hat{y}_{i},\hat{e}_{i}\leftarrow R(L_{i})$triangleright apply regressor head
5:   if $\mathrm{Var}[\hat{y}_{i-2:i}]{<}\tau$ then return $\hat{y}_{i}$triangleright early-exit
@@ Appendix E Upper Confidence Bound Bandit Ablations
We now describe two variations to the Upper Confidence Bound Bandit algorithm described in Section 4: (1) starting at later layers, and (2) balancing exploration and exploitation.
@@
Given the significant jump in COMET score accuracy within the first few layers of the COMET model (Section 3.1), we decided to explore initializing the algorithm with different starting layers. This carries a higher initial cost as we have to run the first few layers of the COMET model for all candidates but could lead to better-informed exploration with the remaining budget. The results in Section 3.1 show that this leads to only marginal improvements.
@@
The heart of the multi-armed bandit problem is the exploration-exploitation trade-off. In our algorithm this trade-off is controlled by the hyperparameter $\gamma$. The higher one chooses $\gamma$, the higher the Upper Confidence Bound scores for uncertain candidates will be, and therefore the more likely the algorithm will be to explore many candidates. On the other hand, a low $\gamma$ will lead the algorithm to go deep with the most promising candidates, i.e., those with the highest estimated scores. We provide results for two different values for $\gamma$ in Section 3.1, which shows that the default choice is likely the most apt.
@@
At the start (budget, $B=5\%$), only the first layer COMET scores are explored for (almost) all candidates, and by the end (budget, $B=100\%$) the full COMET scores are calculated for all candidates. However, the distributions in between reveal that the bandit often explores the candidates up until certain modal layers. These can be clearly seen (e.g. for budget $B=50\%$) to be layers $3$, $4$, $11$, $18$, and $24$. Looking at Figure 7, it can be seen that these values correspond to layers where the predicted error rates (uncertainties) drop significantly, meaning that the evaluation model is now more certain about these scores relative to scores for other candidates. This explains why the UCB bandit algorithm would then prefer to explore less certain candidates with higher UCB scores.
@@ Appendix F Reranking Results in Tabular Format
Table 4 provides the results from Section 3.1 in tabular format.
@@
Table 4: Quality of the candidates returned by the Upper Confidence Bound bandit. Quality is measured in terms of the average final candidate score and the proportion to top-1 candidates selected. We plot these measures for various evaluation budgets. Cost (or budget) is given relative to calculating the full COMET scores for all candidates ($100\%$). Visualized in Section 3.1.
@@ Appendix G Partial COMET
In this section, we introduce Partial COMET (addressing non-standard input obstacle), ensuring reliable quality assessments for incomplete outputs. We show that also beyond evaluation, partial quality estimation can assess translation quality early in the translation generation process, and help discard unpromising candidates from beam search or sampling-based methods, allowing only the most promising candidates to be further expanded. (Appendix G)
@@ G.1 Modeling
This way, the quality estimation model predicts final translation scores based on just the translation prefix. See Table 6.
@@ G.2 Results
We show the results, as measured by Pearson correlation across WMT23 in Table 7. The full COMET strongly underperforms on partial generations. Therefore, this quality estimation model is not suitable for evaluating incomplete segments during generation. In contrast, the partial quality estimation model correlates much more. This can also be due to human annotators overfocusing on the beginning of the translations when providing assessment scores (Magooda and Marcjan, 2020), or the quality estimation system picking up spurious correlations that give away the score (Zouhar et al., 2024).
@@
We try all combinations of pruning bottom 0%, 25%, 50%, or 75% at each of 25%, 50%, and 75% target lengths. The results in Section 3.1 show that by using Partial COMET, we are more likely to prune lower-quality candidates than with the original COMET, which itself performs better than random pruning.